# What Do You See? Enhancing Zero-Shot Image Classification with Multimodal Large Language Models

### v1

@@ 4.2 Ablation Studies
Table 2: Ablation study on the impact of features used by our method on the classification accuracy. DF refers to the description feature, PF refers to the prediction feature, and IF refers to the image feature. In all datasets, we employed the best class feature as indicated in Table 1. Specifically, we utilized the combined class feature for all datasets except for Caltech-101 [25], where we opted for the class description feature. The best results are highlighted in bold.
@@
From the results in Table 2, it is clear that incorporating all three features (DF, PF, and IF) yields the best performance across most datasets, except for the CIFAR datasets [26]. This discrepancy may arise from the low resolution of CIFAR images (originally 32$\times$32), where utilizing the IF may degrade accuracy compared to using only DF and PF.
Table 3: Ablation study on various fusion approaches using 5,000 images randomly selected from the ImageNet dataset [4]. The best results are highlighted in bold.
CLIP model [3]
@@
Table 3 shows the results of our second set of ablation studies, where we report the results for 5,000 randomly selected images from the ImageNet dataset [4]. We explore the use of different cross-modal embedding models (CLIP [ViT-L/14], CLIP [ViT-B/32], and CLIP [ViT-B/16]) [3], and additionally investigate different fusion approaches. Rather than using the mean feature vector of our input features (DF, PF, and IF), we calculated the similarity between each input feature separately and the dataset class label features. Subsequently, we fused the similarity scores to generate a single similarity score for each class label in each dataset. We explored two fusion methods: averaging and taking the maximum for each class label. As shown, averaging our three features (IF, DF, PF) yields the best results.
@@ 6 Limitations
Our method introduces a new approach by leveraging multimodal LLMs to enhance the accuracy of zero-shot image classification. However, it is important to acknowledge that there are still some limitations inherent in our proposed method. Since our method relies on multiple queries to a multimodal LLM to generate the required features (i.e., DF and PF), there may be potential constraints when running on devices with limited computational power, and it may consume more time compared to other methods. Nevertheless, we believe that advancements in LLMs will lead to models that can run efficiently on lower computational power. This would enable broader accessibility and applicability of such models, such as Gemini [5], GPT [2], and LLaMA [40].
@@
Figure 6: Failure examples of our method, where the initial prediction (and the image description in the first example) adversely influenced our final decision. Results are shown for the Caltech-101 dataset [25] (first row) and the Pets dataset [17] (second row).
@@ Appendix A Additional Details
In the main paper, we presented our method for zero-shot image classification. The inference process of our method is concisely described in Algorithm 1. As part of our method, we employed a set of prompts. Table 4 shows the prompts used for each step discussed in the main paper that employs the LLM (Gemini Pro [5]). Specifically, we detail the prompts used to: 1) conduct zero-shot image classification with Gemini Pro [5], 2) describe a given testing image, and 3) generate class labels descriptions.
The class descriptions were generated using five prompts, as shown in Table 4, with 10 responses generated for each prompt and class, resulting in 50 class descriptions per class label. To encourage diversity in Gemini Pro’s responses, we set the temperature parameter to a high value of 0.99, as done in [1]. An example of implementing our method, including both classifier construction and the inference process, is shown in Code 1.
Table 4: Details of prompts utilized in our work. Each row represents one query task to the LLM. For instance, ‘image classification’ indicates the utilization of LLM to conduct initial zero-shot image classification, which serves as one of the features in our method. The {classes} variable refers to the class labels of the dataset. The {class_label} variable denotes one of the class labels in the given dataset.
@@
Table 5: Additional results on 5,000 images from the ImageNet dataset [4]. Best result is highlighted in yellow
@@ Appendix B Additional Results
In this section, we provide supplementary results to those presented in the main paper. Figure 7 shows the confusion matrix for CLIP (ViT-L/14) [3] and our method across two datasets (Caltech-101 [25] and CIFAR-100 [26]). The shown results demonstrate that our method enhances classification accuracy and reduces misclassification rates.
Figure 7: Confusion matrices for zero-shot image classification results of (A) CLIP (ViT-L/14) [3] and (B) our method on the Caltech-101 [25] and CIFAR-100 [26] datasets.
Figure 8: Additional examples demonstrating the influence of input data on final predictions. Examples are provided from the following datasets: SUN397 [21, 22] (first row), DTD [24] (second row), and ImageNet [4] (last three rows).
Table 6: Top-5 classification accuracy of CLIP [3], CuPL [1], and our method on the following datasets: ImageNet [4], CIFAR-10 (C-10) [26], CIFAR-100 (C-100) [26], Food-101 [20], SUN397 [21, 22], Cars [23], DTD [24], Caltech-101 [25], Pets [17], and Places [19]. We report our results with the following class label features: 1) class descriptions, 2) class labels, 3) the template “A photo of {class}”, and 4) combined features of (1-3). The best results are highlighted in yellow.
@@
CLIP (ViT-L/14) [3] 88.4 98.5 77.0 97.8 89.1 93.7 72.8 95.2 96.6 64.5
CuPL [1] 91.0 98.1 79.3 98.3 92.1 94.2 77.7 99.8 96.2 68.8
@@
Table 7: Cohen’s Kappa score of CLIP [3], CuPL [1], and our method on the following datasets: ImageNet [4], CIFAR-10 (C-10) [26], CIFAR-100 (C-100) [26], Food-101 [20], SUN397 [21, 22], Cars [23], DTD [24], Caltech-101 [25], Pets [17], and Places [19]. We report our results with the following class label features: 1) class descriptions, 2) class labels, 3) the template “A photo of {class}”, and 4) combined features of (1-3). The best results are highlighted in yellow.
@@
CLIP (ViT-L/14) [3] 0.651 0.862 0.539 0.867 0.611 0.656 0.452 0.830 0.835 0.372
CuPL [1] 0.665 0.851 0.573 0.889 0.652 0.637 0.480 0.902 0.795 0.395
@@
In Table 2, encouraging results were demonstrated by utilizing the feature of initial prediction produced by the LLM (i.e., Gemini Pro [5]) for zero-shot image classification. Based on these results, one might argue for the direct utilization of Gemini’s class prediction, aiming to match a specific class label from the dataset. However, in several cases, Gemini’s response does not precisely match one of the class labels (as shown in Figure 5). For example, if a ground-truth class label is ‘cat’, Gemini’s response might be ‘The image class is cat’. This discrepancy motivated us to report results of using only Gemini prediction.
In this section, we present additional results from early experiments aimed at utilizing Gemini’s predictions to precisely match one of the class labels in the given dataset. Specifically, we randomly selected 5,000 images from ImageNet [4] for evaluation. While our method, as presented in the main paper, offers a practical way of utilizing Gemini’s predictions, we also present the results of some alternative approaches aimed at precisely identifying one of the dataset class labels, rather than solely relying on the class prediction text generated by Gemini.
Table 5 show the results on the 5,000 images from ImageNet [4] of our main method and alternatives that utilize Gemini’s class prediction to conduct similarity matching with the target dataset class labels. Specifically, we report the results of encoding Gemini’s class prediction using an open-vocabulary language model and measuring the similarity with the encoded class label features. Here, we show the results of using CLIP (ViT-L/14, ViT-B/32, ViT-B/16) [3], DistilBERT [41], and RoBERTa [42].
In addition, we explored classical text similarity metrics – namely, ROUGE-N-F1 and ROUGE-F1 [43] – rather than encoding both Gemini’s prediction and class labels using an open-vocabulary encoding model. As shown in Table 5, our method, which utilizes Gemini’s class prediction as one of the input features, achieves the best results when compared with the alternative approaches.
In the main paper, we reported the top-1 classification accuracy on several datasets [4, 25, 20, 23, 26, 24, 19, 21, 22, 17]. Table 6 presents the top-5 classification accuracy of our method compared to prior work, while Table 7 shows the Cohen’s kappa coefficient. As can be seen, our method achieves a notable improvement while remaining simple and easy to implement.

### v2

@@ 3.2 Ablation Studies
Table 2: Ablation study on the impact of features used by our method on the classification accuracy. DF refers to the description feature, PF refers to the prediction feature, and IF refers to the image feature. In all datasets, we employed the best class feature as indicated in Table 1. Specifically, we utilized the combined class feature for all datasets except for Caltech-101 (Fei-Fei et al. (2004)), where we opted for the class description feature. The best results are highlighted in bold.
@@
From the results in Table 2, it is clear that incorporating all three features (DF, PF, and IF) yields the best performance across most datasets, except for the CIFAR datasets (Krizhevsky et al. (2009)). This discrepancy may arise from the low resolution of CIFAR images (originally 32$\times$32), where utilizing the IF may degrade accuracy compared to using only DF and PF.
Table 3: Ablation study on various fusion approaches using 5,000 images randomly selected from the ImageNet dataset (Deng et al. (2009)). The best results are highlighted in bold.
CLIP model (Radford et al. (2021))
@@
Table 3 shows the results of our second set of ablation studies, where we report the results for 5,000 randomly selected images from the ImageNet dataset (Deng et al. (2009)). We explore the use of different cross-modal embedding models (CLIP [ViT-L/14], CLIP [ViT-B/32], and CLIP [ViT-B/16]) (Radford et al. (2021)), and additionally investigate different fusion approaches. Rather than using the mean feature vector of our input features (DF, PF, and IF), we calculated the similarity between each input feature separately and the dataset class label features. Subsequently, we fused the similarity scores to generate a single similarity score for each class label in each dataset. We explored two fusion methods: averaging and taking the maximum for each class label. As shown, averaging our three features (IF, DF, PF) yields the best results.
@@ 5 Limitations
Our method introduces a new approach by leveraging multimodal LLMs to enhance the accuracy of zero-shot image classification. However, it is important to acknowledge that there are still some limitations inherent in our proposed method. Since our method relies on multiple queries to a multimodal LLM to generate the required features (i.e., DF and PF), there may be potential constraints when running on devices with limited computational power, and it may consume more time compared to other methods. Nevertheless, we believe that advancements in LLMs will lead to models that can run efficiently on lower computational power. This would enable broader accessibility and applicability of such models, such as Gemini (Gemini Team Google (2023)), GPT (Brown et al. (2020)), and LLaMA (Touvron et al. (2023)).
@@
Figure 6: Failure examples of our method, where the initial prediction (and the image description in the first example) adversely influenced our final decision. Results are shown for the Caltech-101 dataset (Fei-Fei et al. (2004)) (first row) and the Pets dataset (Parkhi et al. (2012)) (second row).
@@ Appendix A Additional Details
In the main paper, we presented our method for zero-shot image classification. The inference process of our method is concisely described in Algorithm 1. As part of our method, we employed a set of prompts. Table 4 shows the prompts used for each step discussed in the main paper that employs the LLM (Gemini Pro (Gemini Team Google (2023))). Specifically, we detail the prompts used to: 1) conduct zero-shot image classification with Gemini Pro (Gemini Team Google (2023)), 2) describe a given testing image, and 3) generate class labels descriptions.
The class descriptions were generated using five prompts, as shown in Table 4, with 10 responses generated for each prompt and class, resulting in 50 class descriptions per class label. To encourage diversity in Gemini Pro’s responses, we set the temperature parameter to a high value of 0.99, as done in Pratt et al. (2023). An example of implementing our method, including both classifier construction and the inference process, is shown in Code 1.
Table 4: Details of prompts utilized in our work. Each row represents one query task to the LLM. For instance, ‘image classification’ indicates the utilization of LLM to conduct initial zero-shot image classification, which serves as one of the features in our method. The {classes} variable refers to the class labels of the dataset. The {predicted_class} refers to Gemini Pro’s output of the image classification prompt. The {class_label} variable denotes one of the class labels in the given dataset.
@@
Table 5: Additional results on 5,000 images from the ImageNet dataset (Deng et al. (2009)). Best result is highlighted in yellow
@@ Appendix B Additional Results
In this section, we provide supplementary results to those presented in the main paper. Figure 7 shows the confusion matrix for CLIP (ViT-L/14) (Radford et al. (2021)) and our method across two datasets (Caltech-101 (Fei-Fei et al. (2004)) and CIFAR-100 (Krizhevsky et al. (2009))). The shown results demonstrate that our method enhances classification accuracy and reduces misclassification rates.
Figure 7: Confusion matrices for zero-shot image classification results of (A) CLIP (ViT-L/14) (Radford et al. (2021)) and (B) our method on the Caltech-101 (Fei-Fei et al. (2004)) and CIFAR-100 (Krizhevsky et al. (2009)) datasets.
Figure 8: Additional examples demonstrating the influence of input data on final predictions. Examples are provided from the following datasets: SUN397 (Xiao et al. (2010; 2016)) (first row), DTD (Cimpoi et al. (2014)) (second row), and ImageNet (Deng et al. (2009)) (last three rows).
Table 6: Top-5 classification accuracy of CLIP (Radford et al. (2021)), CuPL (Pratt et al. (2023)), and our method on the following datasets: ImageNet (Deng et al. (2009)), CIFAR-10 (C-10) (Krizhevsky et al. (2009)), CIFAR-100 (C-100) (Krizhevsky et al. (2009)), Food-101 (Bossard et al. (2014)), SUN397 (Xiao et al. (2010; 2016)), Cars (Krause et al. (2013)), DTD (Cimpoi et al. (2014)), Caltech-101 (Fei-Fei et al. (2004)), Pets (Parkhi et al. (2012)), and Places (Zhou et al. (2017)). We report our results with the following class label features: 1) class descriptions, 2) class labels, 3) the template “A photo of {class}”, and 4) combined features of (1-3). The best results are highlighted in yellow.
@@
CLIP (ViT-L/14) (Radford et al. (2021)) 88.4 98.5 77.0 97.8 89.1 93.7 72.8 95.2 96.6 64.5
CuPL (Pratt et al. (2023)) 91.0 98.1 79.3 98.3 92.1 94.2 77.7 99.8 96.2 68.8
@@
Table 7: Cohen’s Kappa score of CLIP (Radford et al. (2021)), CuPL (Pratt et al. (2023)), and our method on the following datasets: ImageNet (Deng et al. (2009)), CIFAR-10 (C-10) (Krizhevsky et al. (2009)), CIFAR-100 (C-100) (Krizhevsky et al. (2009)), Food-101 (Bossard et al. (2014)), SUN397 (Xiao et al. (2010; 2016)), Cars (Krause et al. (2013)), DTD (Cimpoi et al. (2014)), Caltech-101 (Fei-Fei et al. (2004)), Pets (Parkhi et al. (2012)), and Places (Zhou et al. (2017)). We report our results with the following class label features: 1) class descriptions, 2) class labels, 3) the template “A photo of {class}”, and 4) combined features of (1-3). The best results are highlighted in yellow.
@@
CLIP (ViT-L/14) (Radford et al. (2021)) 0.651 0.862 0.539 0.867 0.611 0.656 0.452 0.830 0.835 0.372
CuPL (Pratt et al. (2023)) 0.665 0.851 0.573 0.889 0.652 0.637 0.480 0.902 0.795 0.395
@@
In Table 2, encouraging results were demonstrated by utilizing the feature of initial prediction produced by the LLM (i.e., Gemini Pro (Gemini Team Google (2023))) for zero-shot image classification. Based on these results, one might argue for the direct utilization of Gemini’s class prediction, aiming to match a specific class label from the dataset. However, in several cases, Gemini’s response does not precisely match one of the class labels (as shown in Figure 5). For example, if a ground-truth class label is ‘cat’, Gemini’s response might be ‘The image class is cat’. This discrepancy motivated us to report results of using only Gemini prediction.
In this section, we present additional results from early experiments aimed at utilizing Gemini’s predictions to precisely match one of the class labels in the given dataset. Specifically, we randomly selected 5,000 images from ImageNet (Deng et al. (2009)) for evaluation. While our method, as presented in the main paper, offers a practical way of utilizing Gemini’s predictions, we also present the results of some alternative approaches aimed at precisely identifying one of the dataset class labels, rather than solely relying on the class prediction text generated by Gemini.
Table 5 show the results on the 5,000 images from ImageNet (Deng et al. (2009)) of our main method and alternatives that utilize Gemini’s class prediction to conduct similarity matching with the target dataset class labels. Specifically, we report the results of encoding Gemini’s class prediction using an open-vocabulary language model and measuring the similarity with the encoded class label features. Here, we show the results of using CLIP (ViT-L/14, ViT-B/32, ViT-B/16) (Radford et al. (2021)), DistilBERT (Sanh et al. (2019)), and RoBERTa (Liu et al. (2019)).
In addition, we explored classical text similarity metrics – namely, ROUGE-N-F1 and ROUGE-F1 (Lin (2004)) – rather than encoding both Gemini’s prediction and class labels using an open-vocabulary encoding model. As shown in Table 5, our method, which utilizes Gemini’s class prediction as one of the input features, achieves the best results when compared with the alternative approaches.
In the main paper, we reported the top-1 classification accuracy on several datasets (Deng et al. (2009); Fei-Fei et al. (2004); Bossard et al. (2014); Krause et al. (2013); Krizhevsky et al. (2009); Cimpoi et al. (2014); Zhou et al. (2017); Xiao et al. (2010; 2016); Parkhi et al. (2012)). Table 6 presents the top-5 classification accuracy of our method compared to prior work, while Table 7 shows the Cohen’s kappa coefficient. As can be seen, our method achieves a notable improvement while remaining simple and easy to implement.